# Evolutionary Prediction Games

### v2

@@ B.2 Long-term Fairness
The long-term fairness literature studies the long-term implications of population dynamics on algorithmic fairness (see [16] for a recent survey). In particular, several works in this area also consider different variations of replicator (and replicator-like) dynamics as a model for user response [e.g., 19, 40, 53, 11]. Perhaps closest to our approach is Hashimoto et al. [19], who study worst-group accuracy dynamics under repeated training in a mixture population that evolves via stochastic replicator-like dynamics supplemented by a strong, balanced influx of new users that join the system regardless of its performance (parameterized by $b_{k}$ in their model). While their stability guarantees rely on the assumption of this strong user flow, our results offer a complementary view, showing that stable coexistence may emerge even without an exogenous user flow, under standard empirical risk minimization, and without compromising worst-group accuracy (see Section 5). We also extend their ERM instability results to a wider family of behavioral dynamics, and establish global convergence guarantees, rather than local (see Section 4, and Appendix C).
@@ B.3 Dynamics-Aware Learners
Finally, our framework models the learning algorithm as an entity which is only informed by the current state of population $h\sim\mathcal{A}({\bm{p}})$. While still common in many practical scenarios (see, e.g., [12, 24, 25]), we also note that various practical systems already optimize for long-term objectives. Examples includes learners that reduce hidden incentives to modify the population of users [28, Sec. 7.3], adaption to shifting user preferences [23, 4], or directly maximizing long-term engagement via reinforcement learning (e.g., [1]). While our focus here is on local learning rules, analyzing dynamics-aware learners from an evolutionary perspective is an intriguing direction for future inquiry.
@@ Appendix C Microfoundations
Natural selection dynamics appear in a variety of settings. This appendix offers a representative (though not exhaustive) set of examples showing how local, myopic user behaviors aggregate into population-level dynamics that satisfy our core assumptions. Using online social platforms as a running example, we show how different behavioral patterns (e.g., invitation, imitation, resource allocation) and different types group affiliations (e.g., social, behavioral, demographic) may give rise to natural-selection dynamics in various parts of the ecosystem. We note that similar microfoundations can be identified across a wide range of domains, such as finance, healthcare, and education.
@@ C.2 Imitation
At each step, a content creator observes another creator’s content (e.g., on social media). If their own strategy is less successful than the one they observe, they adopt the other’s strategy with probability proportional to the difference in utilities. Denote the total size of the population by $N$ (here assumed to be large but fixed), the number of creators who adopt strategy $k$ by $N_{k}$, and the proportion of group $k$ by $p_{k}=N_{k}/N$. Sandholm [43, Chapter 4] shows that this form of pairwise proportional imitation converges in the large-population limit to the replicator dynamics [48], which satisfy our core assumptions.
@@ C.3 Online resource allocation
Natural-selection dynamics between groups of users can also arise from interaction between prediction algorithms in multi-stage pipelines. Consider a firm that acquires users though $K$ marketing campaigns, each targeting a different demographic. The firm uses a classifier to personalize the experience of incoming users, and each accurate prediction yields a conversion with probability $\alpha$. Campaign $k$ therefore generates expected utility:
@@
where $\beta>1$ is a hyper-parameter of the MW algorithm. As $\beta\to 1$, dynamics converge towards the replicator dynamics (cf. Equation 17), which satisfies our core assumptions.
@@ D.1 Preliminaries
Figure 5: Graphical representation of two-group evolutionary prediction games (Definition 1). Groups are denoted by ${\left\{{A,B}\right\}}$. Solid lines describe expected group accuracies for each population state ${\bm{p}}$, dotted line describes population accuracy $\mathrm{acc}_{\bm{p}}(h)$, dots indicate Nash equilibria, and x-axis arrows indicate evolutionary stability. Schematic plots illustrate a selection of possible scenarios: (Left) Competitive exclusion. There are two stable equilibria, each dominated by a single group, with locally maximal population accuracy. Additionally, there is an unstable coexistence equilibrium with lower overall accuracy. (Center Left) Stable mutualistic coexistence. There is a stable coexistence equilibrium, where the accuracy of both groups is maximized. (Center Right) Unstable mutualisitic coexistence. Coexistence has maximal population accuracy, but it is unstable. (Right) No coexistence equilibria. There is a single equilibrium dominated by a single group.
@@
We follow the notations of Sandholm [43]. Population games associate each group $k\in[K]$ with a fitness function $F_{k}({\bm{p}})$, which maps a population state ${\bm{p}}\in\Delta^{K}$ to the evolutionary fitness of group $k$. A Nash Equilibrium (NE) of a population game is a population state ${\bm{p}}^{*}$ which satisfies $\operatorname*{support}({\bm{p}}^{*})\subseteq\operatorname*{arg\,max}_{k\in[K]}F_{k}({\bm{p}}^{*})$, and every population game admits at least one Nash equilibrium [43, Theorem 2.1.1]. A population game is a potential game if it can be represented as a gradient of some function. Formally, $F:\Delta^{K}\to\mathbb{R}^{K}$ is a potential game if there exists $f:\Delta^{K}\to\mathbb{R}$ such that $F=\nabla f$, where $\nabla f$ is the gradient of $f$ over the simplex (see [43], Section 3.2 for rigorous definitions).
@@
Coexistence equilibria are fixed points in which two groups or more persist ${\left|{\operatorname*{support}({\bm{p}}^{*})}\right|}\geq 2$. We say that coexistence is competitive if each group’s accuracy at ${\bm{p}}^{*}$ is lower than its accuracy in isolation (i.e., when it is the sole group in the population). Conversely, we say that interactions are mutualistic if every group’s accuracy at ${\bm{p}}^{*}$ exceeds the accuracy it obtains in isolation, indicating that coexistence benefits all groups (see Figure 5 for an illustration). Additionally, we say an equilibrium is evolutionarily stable if it attracts nearby population states (see [43, Section 8.3] for formal definitions).
@@ D.2 Evolutionary Dynamics
For our theoretical analysis, we assume that ${\bm{p}}$ evolves in time according to an ordinary differential equation of the form $\dot{{\bm{p}}}=V_{F}({\bm{p}})$, where $V_{F}:\Delta^{K}\to T\Delta^{K}$ is a vector field tangent to the simplex, induced by the fitness function $F({\bm{p}})$, assumed to be continuous (see [43, Theorem 4.4.1]). A common example for such dynamical system is the replicator equation $\dot{p}_{k}=p_{k}(F_{k}({\bm{p}})-\sum_{k^{\prime}}F_{k^{\prime}}({\bm{p}}))$, which corresponds to the tangent vector field $V_{F}({\bm{p}})={\bm{p}}\odot\left(F({\bm{p}})-\sum_{k^{\prime}}p_{k^{\prime}}F_{k^{\prime}}({\bm{p}}){\bm{1}}\right)$, where $\odot$ is the Hadamard product (see Figure 6). The replicator dynamics are a common model of evolutionary dynamics (see Appendix C, and e.g. [49, 26]), and satisfies our core assumptions. Moreover, we also emphasize that our analysis applies to any $V_{F}$ which satisfies the core assumptions below.
Assumptions.
@@
  • Positive correlation: We assume that $V_{F}({\bm{p}})$ is directionally aligned with the fitness vector $F({\bm{p}})$ whenever a population is not at rest, formally $V_{F}({\bm{p}})\cdot F({\bm{p}})>0$ whenever $V_{F}({\bm{p}})\neq{\bm{0}}$.
  • Nash stationarity: We assume that fixed points $V_{F}({\bm{p}}^{*})={\bm{0}}$ correspond to Nash equilibria ${\bm{p}}^{*}$ of the game $F$ (see eq. 3 for the definition of Nash equilibrium).
For a detailed discussion of the positive correlation assumption, see [43, Section 5.2]. We also note that imitative dynamics (such as the replicator equation), may admit additional rest points at restricted equilibria of the population game (formally, states satisfying $F_{k}({\bm{p}}^{*})=F_{k^{\prime}}({\bm{p}}^{*})$ for all $k,k^{\prime}\in\operatorname*{support}({\bm{p}}^{*})$). However, we note that this does not affect our main results (Theorem 1 and its extensions, and our constructions in Section 5), as any Nash equilibrium is also a restricted equilibrium, and any non-Nash restricted equilibrium corresponds to an unstable fixed point of the imitative dynamics (see [43, Section 8.1]). See Figure 6 for an illustration of Nash and restricted equilibria in a three-group setting.
@@
Dynamics induced by potential games are closely related to the structure of the potential function $f({\bm{p}})$. We informally state key results that are relevant to our setting: By [43, Theorem 3.1.3], a state ${\bm{p}}^{*}$ is a Nash equilibrium of $F$ if and only if it is an extremum of the potential function $f$. By [43, Lemma 7.1.1] potential function are Lyapunov functions, and by [43, Theorem 7.1.2], dynamics satisfying the positive correlation property converge towards Nash equilibria (restricted Nash equilibria in case of replicator dynamics). Finally, by [43, Theorem 8.2.1], a population state ${\bm{p}}^{*}$ is a stable equilbrium if and only if it is a local maximizer of $f({\bm{p}})$.
@@ D.3.1 Linearity
Proposition 1 (Expected loss is linear in ${\bm{p}}$ for a fixed hypothesis).
@@ D.3.2 Overall Accuracy Equality
Equilibria of evolutionary prediction games also satisfy a natural fairness criterion. A classifier $h$ satisfies overall accuracy equality if all groups have equal prediction accuracy. In the context of natural selection, we consider this with respect to the groups currently present in the population:
Definition 1 (Overall accuracy equality; [e.g., 51]).
@@
Definition 2 (Overall accuracy equality in expectation).
@@ D.4 Training Once
Definition 3 (Train-once prediction game).
@@
Definition 4 (Constant game; [43, Section 3.2.4]).
@@
Fix $h_{0}\in\mathcal{Y}^{\mathcal{X}}$. In the train-once setting, the expected marginal accuracies are constant as a function of ${\bm{p}}$, and therefore $F_{k}$ is a constant game by definition 4. By [43, Proposition 3.2.13], $F_{k}$ a potential game with linear potential $f({\bm{p}})=-\vec{\mathcal{L}}(h^{0})\cdot{\bm{p}}$. By [43, Theorem 3.1.3], the set of Nash equilibria is the set of local maximizers of $f({\bm{p}})$, which is a convex combination over the set $\operatorname*{arg\,min}_{k\in[K]}\mathcal{L}_{k}(h^{0})$ due to linearity. Hence:
@@
At equilibrium ${\bm{p}}^{*}$, it holds that $\mathcal{L}_{{\bm{p}}^{*}}=\min_{k\in[K]}\mathcal{L}_{k}(h^{0})$. Apply proposition 1 to obtain:
@@
Finally, observe that each ${\bm{p}}^{*}$ is stable as a maximizer of the linear potential function [43, Theorem 8.2.1]. ∎
Proof of Theorem 1.
Theorem 1 is a special case of Proposition 2 for the loss function: $\mathcal{L}_{k}(h)=-\mathrm{acc}_{k}(h)$. ∎
@@ D.5 Retraining
Figure 7: Illustration of the proof of Theorem 1. Pink line: The expected accuracy of a fixed classifier $h$ is linear in ${\bm{p}}$ (Proposition 1). Blue line: The overall accuracy of the oracle classifier $h_{\bm{p}}$ is a pointwise maximum over linear functions, and is therefore convex (Proposition 3). The gradient of $\mathrm{acc}_{\bm{p}}(h_{\bm{p}})$ over the simplex is the vector of marginal accuracies, showing that the overall accuracy is a potential function for the game (Proposition 4). Properties of the game’s equilibria then follow from convexity arguments.
Definition 5 (Retraining prediction game with an oracle predictor).
@@
Definition 6 (Optimal loss function $l^{*}({\bm{p}})$).
@@
From proposition 1, each $\mathcal{L}_{\bm{p}}(h)$ is linear in ${\bm{p}}$ for all $h\in\mathcal{H}$, and in particular concave. Hence, $l^{*}({\bm{p}})$ is a point-wise minimum of concave functions, and is therefore concave. ∎
Definition 7 (Subgradient; [e.g., 42, Section 23]).
@@
Proposition 4.
Let $h_{\bm{p}}\in\operatorname*{arg\,min}_{h\in\mathcal{H}}\mathcal{L}_{\bm{p}}(h)$ be an optimal predictor. The projection of the loss vector of $h$ at point ${\bm{p}}$ is a subgradient of $-l^{*}({\bm{p}})$:
@@
Let $h_{\bm{p}}\in\operatorname*{arg\,min}_{h\in\mathcal{H}}\mathcal{L}_{\bm{p}}(h)$, and let ${\bm{p}},{\bm{p}}^{\prime}\in\Delta^{K}$. We need to show that $-l^{*}({\bm{p}})+\vec{\mathcal{L}}(h_{\bm{p}})\cdot({\bm{p}}^{\prime}-{\bm{p}})\leq-l^{*}({\bm{p}}^{\prime})$.
By proposition 1, it holds that $\mathcal{L}_{{\bm{p}}^{\prime}}(h_{\bm{p}})=\sum_{k}p^{\prime}_{k}\mathcal{L}_{k}(h_{\bm{p}})$. By algebraic manipulation we obtain:
@@
By proposition 3, it holds that $\mathcal{L}_{{\bm{p}}^{\prime}}(h_{\bm{p}})\geq l^{*}({\bm{p}}^{\prime})$, and therefore:
@@
$l^{*}({\bm{p}})$ is concave, and therefore its negation is convex. Multiplying both sides by $-1$, we obtain that $-{\bm{\Phi}}\vec{\mathcal{L}}({\bm{p}})$ is a subgradient of $-l^{*}$ at point ${\bm{p}}$, as required. ∎
Proposition 5.
Assume $l^{*}({\bm{p}})$ is differentiable, and let $h_{\bm{p}}\in\operatorname*{arg\,min}_{h\in\mathcal{H}}\mathcal{L}_{\bm{p}}(h)$. The gradient of $-l^{*}({\bm{p}})$ satisfies:
@@
By proposition 4, it holds that $-{\bm{\Phi}}\vec{\mathcal{L}}(h_{\bm{p}})\in\partial\left(-l^{*}(p)\right)$, and therefore $\nabla\left(-l^{*}({\bm{p}})\right)=-{\bm{\Phi}}\vec{\mathcal{L}}(h_{\bm{p}})$. ∎
Lemma 1.
@@
Take the function $f({\bm{p}})=-l^{*}({\bm{p}})$. By proposition 5, the gradient of $f({\bm{p}})$ satisfies:
@@
Definition 9 (Oracle-equivalent groups).
@@
If the groups are oracle-equivalent, then by Definition 9 there exists
@@
By Proposition 3, the optimal loss $l^{*}$ is concave over the simplex $\Delta^{K}$, and therefore its minimizers are convex combinations of simplex vertices for which the function is constant. By Proposition 6, such combinations correspond to sets of oracle-equivalent groups. ∎
Proposition 8.
@@
When each group has a distinct optimal classifier, the only sets of oracle-equivalent groups are the singleton sets ${\left\{{1}\right\}},\dots,{\left\{{K}\right\}}$. The claim then follows from Proposition 7. ∎
Proof of Theorem 1.
By applying Lemma 1 on the loss function $\mathcal{L}_{k}(h)=-\mathrm{acc}_{k}(h)$, we obtain that $F({\bm{p}})$ is a potential game with potential function $f({\bm{p}})=\mathrm{acc}_{\bm{p}}(h_{\bm{p}})$. By Proposition 3, we obtain that $f({\bm{p}})$ is convex over the simplex. When no two groups share an optimal classifier Proposition 8 holds, and local maximizers of $f$ exist and are located at the vertices of the simplex. By [43, Theorem 8.2.1], maximizers of $f({\bm{p}})$ correspond to evolutionarily stable Nash equilibria of $F({\bm{p}})$, and therefore any stable equilibrium ${\bm{p}}^{*}$ satisfies ${\left|{\operatorname*{support}({\bm{p}}^{*})}\right|}=1$. Additionally, we note that coexistence equilibria may exist as $f({\bm{p}})$ may have local minimizers. This proves claims 2,3,4 (Stability, Exclusion, Coexistence) in the statement of the theorem. Finally, for claim 1 (Accuracy), by [43, Lemma 7.1.1] the potential function $f({\bm{p}})$ is a Lyapunov function for any evolutionary dynamic which satisfies the positive correlation property, and therefore $\tfrac{\mathrm{d}}{\mathrm{d}t}f({\bm{p}})=\tfrac{\mathrm{d}}{\mathrm{d}t}\mathrm{acc}_{{\bm{p}}}(h_{{\bm{p}}})\geq 0$. ∎

### v3

@@ B.2 Long-Term Fairness
Group fairness is often formalized through measures of group disparity, such as the difference in error rates across groups (see, e.g., [18, 2]). However, these measures may fail to capture fairness in dynamic settings. The long-term fairness literature therefore studies the long-term implications of population dynamics on algorithmic fairness; one straightforward example is that if a classifier is trained to comply with fairness constraints, dynamics can cause this guarantee to erode over time [40] (See [22] for a recent survey). Our work proposes a complementing (and somewhat nuanced) counterfactual perspective, suggesting that a system can appear to comply with a fairness definition, but only because some groups were already driven out of the system, and disparity is measured only over the groups that remain.
Several works in the long-term fairness literature also consider replicator (and replicator-like) dynamics of different kinds and variations as models for user response behavior [e.g., 25, 49, 63, 15], albeit not in an evolutionary game setting. Perhaps closest to our approach is Hashimoto et al. [25], who study worst-group accuracy dynamics under repeated training in a mixture population that evolves via stochastic replicator-like dynamics supplemented by a strong, balanced influx of new users that join the system regardless of its performance (parameterized by $b_{k}$ in their model). While their stability guarantees rely on the assumption of this strong user flow, our results offer a complementary view, showing that stable coexistence may emerge even without an exogenous user flow, under standard empirical risk minimization, and without compromising worst-group accuracy (see Section˜5). We also extend their ERM instability results to a wider family of behavioral dynamics, and establish global convergence guarantees, rather than local (see Section˜4, and Appendix˜C).
@@ B.3 Dynamics-Aware Learners
Finally, our framework models the learning algorithm as an entity which is only informed by the current state of population $h\sim\mathcal{A}({\bm{p}})$, and in particular does not take into account the implications of classifier deployment. While still common in many practical scenarios (see, e.g., [16, 33, 34]), we also note that various practical systems already optimize for long-term objectives. Examples includes learners that reduce hidden incentives to modify the population of users [37, Sec. 7.3], adaption to shifting user preferences [31, 6], or directly maximizing long-term engagement via reinforcement learning (e.g., [1]). While our focus here is on local learning rules, analyzing dynamics-aware learners from an evolutionary perspective is an intriguing direction for future inquiry.
@@ Appendix C Microfoundations
Natural selection dynamics appear in a variety of settings. This appendix offers a representative (though not exhaustive) set of examples showing how local, myopic user behaviors aggregate into population‐level dynamics that satisfy our core assumptions. Using online social platforms as a running example, we show how different behavioral patterns (e.g., invitation, imitation, resource allocation) and different types group affiliations (e.g., social, behavioral, demographic) may give rise to natural‐selection dynamics in various parts of the ecosystem. We note that similar microfoundations can be identified across a wide range of domains, such as finance, healthcare, and education.
@@ C.2 Imitation
At each step, a content creator observes another creator’s content (e.g., on social media). If their own strategy is less successful than the one they observe, they adopt the other’s strategy with probability proportional to the difference in utilities. Denote the total size of the population by $N$ (here assumed to be large but fixed), the number of creators who adopt strategy $k$ by $N_{k}$, and the proportion of group $k$ by $p_{k}=N_{k}/N$. Sandholm [52, Chapter 4] shows that this form of pairwise proportional imitation converges in the large-population limit to the replicator dynamics [57], which satisfy our core assumptions.
@@ C.3 Online resource allocation
Natural‐selection dynamics between groups of users can also arise from interaction between prediction algorithms in multi‐stage pipelines. Consider a firm that acquires users though $K$ marketing campaigns, each targeting a different demographic. The firm uses a classifier to personalize the experience of incoming users, and each accurate prediction yields a conversion with probability $\alpha$. Campaign $k$ therefore generates expected utility:
@@
where $\beta>1$ is a hyper-parameter of the MW algorithm. As $\beta\to 1$, dynamics converge towards the replicator dynamics (cf. Equation˜17), which satisfies our core assumptions.
@@ D.1 Preliminaries
Figure 6: Graphical representation of two-group evolutionary prediction games (Definition˜1). Groups are denoted by $\left\{{A,B}\right\}$}. Solid lines describe expected group accuracies for each population state $\bm{p}$}, dotted line describes population accuracy $\mathrm{acc}_{\bm{p}}(h)$, dots indicate Nash equilibria, and x-axis arrows indicate evolutionary stability. Schematic plots illustrate a selection of possible scenarios: (Left) Competitive exclusion. There are two stable equilibria, each dominated by a single group, with locally maximal population accuracy. Additionally, there is an unstable coexistence equilibrium with lower overall accuracy. (Center Left) Stable mutualistic coexistence. There is a stable coexistence equilibrium, where the accuracy of both groups is maximized. (Center Right) Unstable mutualisitic coexistence. Coexistence has maximal population accuracy, but it is unstable. (Right) No coexistence equilibria. There is a single equilibrium dominated by a single group.
@@
We follow the notations of Sandholm [52], and include the key definitions here for completeness. Population games associate each group $k\in[K]$ with a fitness function $F_{k}({\bm{p}})$, which maps a population state ${\bm{p}}\in\Delta^{K}$ to the evolutionary fitness of group $k$. A Nash Equilibrium (NE) of a population game is a population state $\bm{p}$$^{*}$ which satisfies $\operatorname*{support}({\bm{p}}^{*})\subseteq\operatorname*{arg\,max}_{k\in[K]}F_{k}({\bm{p}}^{*})$, and every population game admits at least one Nash equilibrium [52, Theorem 2.1.1]. A Restricted Equilibrium (RE) of a population game is a population state $\bm{p}$$^{*}$ which satisfies $\operatorname*{support}({\bm{p}}^{*})\subseteq\operatorname*{arg\,max}_{k\in\operatorname*{support}({\bm{p}}^{*})}F_{k}({\bm{p}}^{*})$, or equivalently $F_{k}({\bm{p}}^{*})=F_{k^{\prime}}({\bm{p}}^{*})$ for all $k,k^{\prime}\in\operatorname*{support}({\bm{p}}^{*})$. A population game is a potential game if it can be represented as a gradient of some function. Formally, $F:\Delta^{K}\to\mathbb{R}^{K}$ is a potential game if there exists $f:\Delta^{K}\to\mathbb{R}$ such that $F=\nabla f$, where $\nabla f$ is the gradient of $f$ over the simplex (see [52], Section 3.2 for rigorous definitions).
@@
Coexistence equilibria are fixed points in which two groups or more persist ${\left|{\operatorname*{support}({\bm{p}}^{*})}\right|}\geq 2$. We say that coexistence is competitive if each group’s accuracy at $\bm{p}$$^{*}$ is lower than its accuracy in isolation (i.e., when it is the sole group in the population). Conversely, we say that interactions are mutualistic if every group’s accuracy at $\bm{p}$$^{*}$ exceeds the accuracy it obtains in isolation, indicating that coexistence benefits all groups (see Figure˜6 for an illustration). Additionally, we say an equilibrium is evolutionarily stable if it attracts nearby population states (see [52, Section 8.3] for formal definitions).
@@ D.2 Evolutionary Dynamics
For our theoretical analysis, we assume that $\bm{p}$} evolves in time according to an ordinary differential equation of the form $\dot{{\bm{p}}}=V_{F}({\bm{p}})$, where $V_{F}:\Delta^{K}\to T\Delta^{K}$ is a vector field tangent to the simplex, induced by the fitness function $F({\bm{p}})$, assumed to be continuous (see [52, Theorem 4.4.1]). A common example for such dynamical system is the replicator equation $\dot{p}_{k}=p_{k}(F_{k}({\bm{p}})-\sum_{k^{\prime}}F_{k^{\prime}}({\bm{p}}))$, which corresponds to the tangent vector field $V_{F}({\bm{p}})={\bm{p}}\odot\left(F({\bm{p}})-\sum_{k^{\prime}}p_{k^{\prime}}F_{k^{\prime}}({\bm{p}}){\bm{1}}\right)$, where $\odot$ is the Hadamard product (see Figure˜7). The replicator dynamics are a common model of evolutionary dynamics (see Appendix˜C, and e.g. [58, 35]), and satisfies our core assumptions. Moreover, we also emphasize that our analysis applies to any $V_{F}$ which satisfies the core assumptions below.
@@
  • Positive correlation: We assume that $V_{F}({\bm{p}})$ is directionally aligned with the fitness vector $F({\bm{p}})$ whenever a population is not at rest, formally $V_{F}({\bm{p}})\cdot F({\bm{p}})>0$ whenever $V_{F}({\bm{p}})\neq{\bm{0}}$. For a detailed discussion, see [52, Section 5.2].
For correspondence to game equilibria, we assume either of the following assumptions:
  • Nash stationarity: Under Nash stationarity, all fixed points $V_{F}({\bm{p}}^{*})={\bm{0}}$ correspond to Nash equilibria $\bm{p}$$^{*}$ of the game $F$. See eq.˜3 for the definition of Nash equilibrium, and [52, Section 5] for a detailed discussion.
  • Imitative dynamics: Under imitative dynamics, agents adopt strategies based on observing and copying the behaviors of others in a population. See [52, Section 5.4] for a formal definition using the revision protocol formalism.
The replicator equation introduced in Section˜3 is a special case of imitative dynamics, which satisfy our assumptions. In contrast to dynamics with Nash stationarity, imitative dynamics admit additional rest points at restricted equilibria of the population game (formally, states satisfying $F_{k}({\bm{p}}^{*})=F_{k^{\prime}}({\bm{p}}^{*})$ for all $k,k^{\prime}\in\operatorname*{support}({\bm{p}}^{*})$). However, we note that this does not affect our main results (Theorem˜1 and its extensions, and our constructions in Section˜5), as any Nash equilibrium is also a restricted equilibrium, and any non-Nash restricted equilibrium corresponds to an unstable fixed point of the imitative dynamics (see [52, Section 8.1]). See Figure˜7 for an illustration of Nash and restricted equilibria in a three-group setting.
@@
Dynamics induced by potential games are closely related to the structure of the potential function $f({\bm{p}})$. We informally state key results that are relevant to our setting: By [52, Theorem 3.1.3], a state $\bm{p}$$^{*}$ is a Nash equilibrium of $F$ if and only if it is an extremum of the potential function $f$. By [52, Lemma 7.1.1] potential function are Lyapunov functions, and by [52, Theorem 7.1.2], dynamics satisfying the positive correlation property converge towards Nash equilibria (restricted Nash equilibria in case of replicator dynamics). Finally, by [52, Theorem 8.2.1], a population state $\bm{p}$$^{*}$ is a stable equilbrium if and only if it is a local maximizer of $f({\bm{p}})$.
@@ D.3.1 Linearity
Proposition D.1 (Expected loss is linear in $\bm{p}$} for a fixed hypothesis).
@@ D.3.2 Overall Accuracy Equality
Equilibria of evolutionary prediction games also satisfy a natural fairness criterion. A classifier $h$ satisfies overall accuracy equality if all groups have equal prediction accuracy [61]. In the context of natural selection, we consider this with respect to the groups currently present in the population:
Definition D.1 (Overall accuracy equality; [e.g., 61]).
@@
Definition D.2 (Overall accuracy equality in expectation).
@@
Furthermore, we note that similar reasoning can be utilized to prove an analogous correspondence between overall accuracy fairness and restricted equilibria of the game.
@@ D.4 Training Once
Definition D.3 (Train-once prediction game).
@@
Definition D.4 (Constant game; [52, Section 3.2.4]).
@@
Fix $h_{0}\in\mathcal{Y}^{\mathcal{X}}$. In the train-once setting, the expected marginal accuracies are constant as a function of $\bm{p}$}, and therefore $F_{k}$ is a constant game by definition˜D.4. By [52, Proposition 3.2.13], $F_{k}$ a potential game with linear potential $f({\bm{p}})=-\vec{\mathcal{L}}(h^{0})\cdot{\bm{p}}$. By [52, Theorem 3.1.3], the set of Nash equilibria is the set of local maximizers of $f({\bm{p}})$, which is a convex combination over the set $\operatorname*{arg\,min}_{k\in[K]}\mathcal{L}_{k}(h^{0})$ due to linearity. Hence:
@@
At equilibrium $\bm{p}$$^{*}$, it holds that $\mathcal{L}_{{\bm{p}}^{*}}=\min_{k\in[K]}\mathcal{L}_{k}(h^{0})$. Apply proposition˜D.1 to obtain:
@@
Finally, observe that each $\bm{p}$$^{*}$ is stable as a maximizer of the linear potential function [52, Theorem 8.2.1]. ∎
Proof of Theorem˜D.1.
Theorem˜D.1 is a special case of Proposition˜D.2 for the loss function: $\mathcal{L}_{k}(h)=-\mathrm{acc}_{k}(h)$. ∎
@@ D.5 Retraining
Figure 8: Illustration of the proof of Theorem˜1. Pink line: The expected accuracy of a fixed classifier $h$ is linear in $\bm{p}$} (Proposition˜D.1). Blue line: The overall accuracy of the oracle classifier $h_{\bm{p}}$ is a pointwise maximum over linear functions, and is therefore convex (Proposition˜D.3). The gradient of $\mathrm{acc}_{\bm{p}}(h_{\bm{p}})$ over the simplex is the vector of marginal accuracies, showing that the overall accuracy is a potential function for the game (Proposition˜D.4). Properties of the game’s equilibria then follow from convexity arguments.
Definition D.5 (Retraining prediction game with an oracle predictor).
@@
Definition D.6 (Optimal loss function $l^{*}({\bm{p}})$).
@@
From proposition˜D.1, each $\mathcal{L}_{\bm{p}}(h)$ is linear in $\bm{p}$} for all $h\in\mathcal{H}$, and in particular concave. Hence, $l^{*}({\bm{p}})$ is a point-wise minimum of concave functions, and is therefore concave. ∎
Definition D.7 (Subgradient; [e.g., 51, Section 23]).
@@
Proposition D.4.
Let $h_{\bm{p}}\in\operatorname*{arg\,min}_{h\in\mathcal{H}}\mathcal{L}_{\bm{p}}(h)$ be an optimal predictor. The projection of the loss vector of $h$ at point $\bm{p}$} is a subgradient of $l^{*}({\bm{p}})$:
@@
Let $h_{\bm{p}}\in\operatorname*{arg\,min}_{h\in\mathcal{H}}\mathcal{L}_{\bm{p}}(h)$, and let ${\bm{p}},{\bm{p}}^{\prime}\in\Delta^{K}$. We need to show that $l^{*}({\bm{p}})+\vec{\mathcal{L}}(h_{\bm{p}})\cdot({\bm{p}}^{\prime}-{\bm{p}})\leq-l^{*}({\bm{p}}^{\prime})$.
By proposition˜D.1, it holds that $\mathcal{L}_{{\bm{p}}^{\prime}}(h_{\bm{p}})=\sum_{k}p^{\prime}_{k}\mathcal{L}_{k}(h_{\bm{p}})$. By algebraic manipulation we obtain:
@@
By proposition˜D.3, it holds that $\mathcal{L}_{{\bm{p}}^{\prime}}(h_{\bm{p}})\geq l^{*}({\bm{p}}^{\prime})$, and therefore:
@@
$l^{*}({\bm{p}})$ is concave, and therefore its negation is convex. Multiplying both sides by 1-1, we obtain that ${\bm{\Phi}}\vec{\mathcal{L}}({\bm{p}})$ is a subgradient of $l^{*}$ at point $\bm{p}$}, as required. ∎
Proposition D.5.
Assume $l^{*}({\bm{p}})$ is differentiable, and let $h_{\bm{p}}\in\operatorname*{arg\,min}_{h\in\mathcal{H}}\mathcal{L}_{\bm{p}}(h)$. The gradient of $l^{*}({\bm{p}})$ satisfies:
@@
By proposition˜D.4, it holds that ${\bm{\Phi}}\vec{\mathcal{L}}(h_{\bm{p}})\in\partial\left(-l^{*}(p)\right)$, and therefore $\nabla\left(-l^{*}({\bm{p}})\right)=-{\bm{\Phi}}\vec{\mathcal{L}}(h_{\bm{p}})$. ∎
Lemma D.1.
@@
Take the function $f({\bm{p}})=-l^{*}({\bm{p}})$. By proposition˜D.5, the gradient of $f({\bm{p}})$ satisfies:
@@
Definition D.9 (Oracle-equivalent groups).
@@
If the groups are oracle-equivalent, then by Definition˜D.9 there exists
@@
By Proposition˜D.3, the optimal loss $l^{*}$ is concave over the simplex $\Delta^{K}$, and therefore its minimizers are convex combinations of simplex vertices for which the function is constant. By Proposition˜D.6, such combinations correspond to sets of oracle-equivalent groups. ∎
Proposition D.8.
@@
When each group has a distinct optimal classifier, the only sets of oracle-equivalent groups are the singleton sets ${\left\{{1}\right\}},\dots,{\left\{{K}\right\}}$. The claim then follows from Proposition˜D.7. ∎
Proof of Theorem˜1.
By applying Lemma˜D.1 on the loss function $\mathcal{L}_{k}(h)=-\mathrm{acc}_{k}(h)$, we obtain that $F({\bm{p}})$ is a potential game with potential function $f({\bm{p}})=\mathrm{acc}_{\bm{p}}(h_{\bm{p}})$. By Proposition˜D.3, we obtain that $f({\bm{p}})$ is convex over the simplex. When no two groups share an optimal classifier Proposition˜D.8 holds, and local maximizers of $f$ exist and are located at the vertices of the simplex. By [52, Theorem 8.2.1], maximizers of $f({\bm{p}})$ correspond to evolutionarily stable Nash equilibria of $F({\bm{p}})$, and therefore any stable equilibrium $\bm{p}$$^{*}$ satisfies ${\left|{\operatorname*{support}({\bm{p}}^{*})}\right|}=1$. Additionally, we note that coexistence equilibria may exist as $f({\bm{p}})$ may have local minimizers. This proves claims 2,3,4 (Stability, Exclusion, Coexistence) in the statement of the theorem. Finally, for claim 1 (Accuracy), by [52, Lemma 7.1.1] the potential function $f({\bm{p}})$ is a Lyapunov function for any evolutionary dynamic which satisfies the positive correlation property, and therefore $\tfrac{\mathrm{d}}{\mathrm{d}t}f({\bm{p}})=\tfrac{\mathrm{d}}{\mathrm{d}t}\mathrm{acc}_{{\bm{p}}}(h_{{\bm{p}}})\geq 0$. ∎